OpenStudy (anonymous):

what is the magnitude of the force necessary to accelerate a 20.0 kg mass from rest to 0.500 m/s in a time of 2.0s

OpenStudy (ajprincess):

First find a using v=u+at Here v-final velocity(in this case 0.500m/s) u-initial velocity(in this case 0m/s) t-time taken(in this case 2.0s) a-acceleration Then find the magnitude of the force using F=ma m-mass a-acceleration

OpenStudy (anonymous):

a=v/t. so .5/2 = 1. f=m*a so f=(20)*(1) = 20 N

OpenStudy (anonymous):

v=u+at v/t = a a = 0.5/2 = 1/4m/s^2 f=ma = 20*1/4 = 5N